OpenStudy (anonymous):

find all the the complex fourth roots of -1-i. you may leave answers in polar form

OpenStudy (phi):

The question implies changing to polar form \[ m e^{i \theta} \] m is the magnitude. for a complex number in cartesian form a + bi \[ m= \sqrt{a^2+b^2} \] the angle (measured counter-clockwise from the positive x-axis) is